### v1

# An End-to-End Model for Photo-Sharing Multi-modal Dialogue Generation

Abstract.
Photo-Sharing Multi-modal dialogue generation requires a dialogue agent not only to generate text responses but also to share photos at the proper moment. Using image text caption as the bridge, a pipeline model integrates an image caption model, a text generation model, and an image generation model to handle this complex multi-modal task. However, representing the images with text captions may loss important visual details and information and cause error propagation in the complex dialogue system. Besides, the pipeline model isolates the three models separately because discrete image text captions hinder end-to-end gradient propagation. We propose the first end-to-end model for photo-sharing multi-modal dialogue generation, which integrates an image perceptron and an image generator with a large language model. The large language model employs the Q-Former to perceive visual images in the input end. For image generation in the output end, we propose a dynamic vocabulary transformation matrix and use straight-through and gumbel-softmax techniques to align the large language model and stable diffusion model and achieve end-to-end gradient propagation. We perform experiments on PhotoChat and DialogCC datasets to evaluate our end-to-end model. Compared with pipeline models, the end-to-end model gains state-of-the-art performances on various metrics of text and image generation. More analysis experiments also verify the effectiveness of the end-to-end model for photo-sharing multi-modal dialogue generation.
Multi-modal dialogue, Photo-Sharing, Large language model, Stable diffusion, End-to-end model
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TOMMccs: Computing methodologies Natural language generationccs: Computing methodologies Discourse, dialogue and pragmaticsccs: Computing methodologies Image representationsccs: Computing methodologies Reconstruction
1. Introduction
Human-machine dialogue system is a crucial research and application area for the artificial intelligence community (Zhang et al., 2020; Roller et al., 2021; Adiwardana et al., 2020; Yi et al., 2024). With the development of vision-language representation learning, multi-modal dialogue has attracted widespread attention (Das et al., 2017; Mostafazadeh et al., 2017; Shuster et al., 2020; Zheng et al., 2022; Shuster et al., 2021). The multi-modal settings in these literatures are limited to adding visual modality to the input, while the output of the dialogue model remains pure textual modality. Similar to the conversation manner of humans on instant messaging softwares (e.g., Facebook, WhatsApp, and WeChat), an engaging multi-modal dialogue agent should not only perceive and comprehend visual content but also generate images. Therefore, researchers have started to focus on photo-sharing multi-modal dialogue in recent years (Lee et al., 2021; Zang et al., 2021; Sun et al., 2022; Feng et al., 2023; Lee et al., 2024; Aboutalebi et al., 2024), which requires the active behavior of sharing photos in multi-turn dialogue for multi-modal chatbots.
In order to build the photo-sharing multi-modal dialogue system, Sun et al. (2022) propose a pipeline model, Divter, which composes three modules: image caption model, text dialogue model (DialogGPT (Zhang et al., 2020)), and image generation model (DALL-E (Ramesh et al., 2021)). The pipeline model utilizes the image text description as the bridge to connect visual and textual modality. Specifically, for the input end, the text dialogue model perceives photos based on the image descriptions from the image caption model. For the output end, the text dialogue model generates text responses and image descriptions, which are converted into photos by the image generation model.
Although the pipeline model can achieve photo-sharing multi-modal dialogue, the system still faces many challenges. First, the text dialogue model may lose visual details in the photos, because the image caption model will significantly compress visual information during the process of generating image descriptions. As a result, it is difficult for the model to produce elaborate utterance replies. Second, the pipeline model suffers from the severe problem of error propagation. Any small mistake in the image descriptions will be propagated and enlarged into the generated text dialogue responses and images. Third, although the pipeline model includes three neural network models, they are independent and cannot achieve end-to-end gradient propagation and training. So the separate training of these models makes it difficult to capture the internal relationships between different modules, thereby hindering a unified end-to-end photo-sharing multi-modal dialogue model. Besides, the cascaded system has redundant functions and parameters, leading to latency and inefficiency in practical deployment.
Recent progress in multi-modal representation alignment (Radford et al., 2021; Zhang et al., 2024; Bordes et al., 2024), large language models (LLMs) (Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023) and diffusion models (Ho et al., 2020; Rombach et al., 2022) offer potential solutions and directions for addressing these issues. For one thing, large language models and diffusion models have achieved remarkable results in both text and image generation tasks. Advanced foundation models can improve the final performance of the photo-sharing multi-modal dialogue system. For another thing, current multi-modal large language models integrate visual encoders with large language models, unifying the representation space of vision and text. This enables text generation models to directly perceive the entire image, rather than relying on image captions that may lose visual details. Furthermore, a single multi-modal model performing both image perception and text generation fully liberates the image caption model.
In this work, we propose an end-to-end photo-sharing multi-modal dialogue model to alleviate the challenges in the pipeline model, resorting to recent multi-modal generation models and end-to-end gradient connection techniques (Straight-Through Estimator (Bengio et al., 2013) and Gumbel-Softmax (Jang et al., 2017; Maddison et al., 2017)). First, we use the large language model (Llama 2 (Touvron et al., 2023)) as the backbone of text dialogue generation. Second, Q-Former in BLIP-2 (Li et al., 2023) connects the vision representation encoder (ViT (Dosovitskiy et al., 2021)) with the large language model, which enables the realization of the end-to-end multi-modal model at the input end. Third, the stable diffusion (Rombach et al., 2022) is engaged in image generation in our photo-sharing dialogue system. Especially, we design a dynamic vocabulary transformation matrix to convert LLM vocab into diffusion model vocab, which integrates Straight-Through Estimator (Bengio et al., 2013) and Gumbel-Softmax (Jang et al., 2017; Maddison et al., 2017) to connect the diffusion model with the large language model and realize the end-to-end gradient connection at the output end. Overall, the vision encoder and the diffusion model equip the large language model with the ability to see and draw images at the same time. Q-Former, dynamic vocabulary transformation matrix and Straight-Through Gumbel-Softmax implement end-to-end gradient propagation and joint training.
We conduct experiments on PhotoChat (Zang et al., 2021) and DialogCC (Lee et al., 2024) datasets to verify the effectiveness of our end-to-end photo-sharing multi-modal dialogue model. We employ many diverse textual and visual metrics to evaluate the performances of text and image generation, including BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), Inception Score (IS) (Salimans et al., 2016) and Frechet Inception Distance (FID) (Heusel et al., 2017). Our end-to-end model acquires improvements on all metrics compared with the conventional pipeline model (Divter) and LLM-based pipeline model.
Figure 1. Photo-sharing Multi-modal dialogue.
In conclusion, the main contributions can be summarized as follows:
  • We propose the first end-to-end photo-sharing multi-modal dialogue model, which jointly trains the image encoder, large language model and diffusion model to acquire the abilities of image perceiving, text generation and image generation.
  • We design a novel dynamic vocabulary transformation matrix to convert image text description from LLM to diffusion model while retaining gradients.
  • Experimental results on diverse metrics verify the effectiveness of our end-to-end model compared with pipeline models.
2. Model
Figure 2. The overview of our end-to-end photo-sharing multi-modal dialogue generation model, which integrates an image perceptron and an image generator with a large language model. The large language model employs the Q-Former to perceive visual images in the input end. For image generation in the output end, we propose a dynamic vocabulary transformation matrix (DVTM) and use straight-through and gumbel-softmax (STGS) techniques to align the large language model and stable diffusion model and achieve end-to-end gradient propagation.
In this section, we start with the task formalization of photo-sharing multi-modal dialogue (§2.1), then introduce our end-to-end model, which consists of three components: text dialogue generator (§2.2), image perception (§2.3) and image generator (§2.4), end with the description of end-to-end training of our multi-modal dialogue model (§2.5).
2.1. Task Formalization
Given a photo-sharing multi-modal dialogue dataset $\mathcal{D}=\{(U_{i},R_{i})\}_{i}^{I}$, where $\forall~{}i\in\{1,2,\dots,I\}$, $U_{i}$ is the dialogue context, $R_{i}$ is the response regarding to $U_{i}$. Dialogue context $U_{i}$ and response $R_{i}$ could include two modalities: textual elements (e.g., utterances) and visual elements (e.g., images). Concretely, for each dialogue context and response, $U_{i}=\{u_{k}\}_{k}^{K}$ and $R_{i}=\{r_{l}\}_{l}^{L}$ are sequences of multi-modal elements including textual utterances and visual images. $K$ and $L$ are the number of elements in $U_{i}$ and $R_{i}$ respectively. Our goal is to learn an end-to-end photo-sharing multi-modal dialogue model $P(R|U)$ supervised by the dataset $\mathcal{D}$. When given the multi-modal dialogue context $U$, the end-to-end model can generate multi-modal response $R$ including texts and images.
2.2. Text Dialogue Generator
In recent years, large language models have demonstrated powerful text generation ability. Besides, they can also acquire visual ability through extra vision modules, becoming multi-modal large language models. Therefore, we adapt the large language model, Llama (Touvron et al., 2023), as the backbone of our end-to-end photo-sharing multi-modal dialogue model. Specifically, Llama is a sequence-to-sequence generation model based on decoder-only transformer architecture, which autoregressively generates the next token according to the input sequence.
[…]
We input the dialogue context $U$ and response $R$ to the text dialogue generator as the source and target sequence for teacher-forcing training. First, the large language model generates the hidden state representation of the current time step $\bm{h}_{i}$, based on the dialogue context $U$ and the target sequence of the first $i-1$ time steps $r_{\textless i}$. Then, a linear projection layer (LPL) converts the dimension of the hidden state vector $\bm{h}_{i}$ into the size of the large language model vocabulary. This operation can be seen as computing the similarity vector $\bm{o}_{i}$ between $\bm{h}_{i}$ and all tokens in the vocabulary. Finally, we apply softmax to gain the token probability distribution $\bm{p}_{i}$ on the vocabulary of the large language model for the current time step:
(1) $\bm{h}_{i}$ $=\text{LLM}(U,r_{\textless i})$
$\bm{o}_{i}$ $=\text{LPL}(\bm{h}_{i})$
$\bm{p}_{i}$ $=\text{SoftMax}(\bm{o}_{i})$
In the end, we minimize the negative log-likelihood of the target sequence $y$ to optimize the text dialogue generator:
(2) $\mathcal{L}^{\text{t}}=-\text{log}~{}P(R|U)=-\text{log}\prod_{i=1}^{|R|}~{}P(y_{i}|U,r_{\textless i})=-\sum_{i=1}^{|R|}\text{log}~{}\bm{p}_{i}^{r_{i}}$
where $\mathcal{L}^{\text{t}}$ is the notation of text generation loss.
For the inference stage, we sample the token from vocabulary probability distribution $\bm{p}$ and generate the response sequence auto-regressively. Concretely, the decoder-only-based model predicts the next token by encoding the dialogue context and the historical token sequence of the generated response. The text dialogue generator repeats the generation process until the generated response reaches the end token or the max sequence length. Particularly, contents between image start token [IMG] and image end token [/IMG] will be inputted to the image generator, guiding the diffusion model to generate appropriate images. Besides, we also feed the image text description and the corresponding vocabulary probability distribution $\bm{p}_{i}$ to the image generator during the training stage. Our novel gradient connection technique maintains gradient propagation between the large language model and the diffusion model, and hence builds an end-to-end model, which will represented in the following subsection 2.4.
2.3. Image Perceptron
For the photo in the input end, the pipeline model bridges text and vision modality by discrete image text description, which interrupts gradient computation and overlooks visual details. To enable the large language model to directly perceive the visual modality information of the entire image, rather than relying on image descriptions of textual modality, image perceptron attempts to connect a visual encoder to the large language model. There are two main approaches connecting visual encoders and large language models by inserting extra vision-language adaption layers: Q-Former (Li et al., 2023) and linear projection (Liu et al., 2023). We select Q-Former for our end-to-end photo-sharing multi-modal dialogue model, because it encodes images with arbitrary size to a fixed number of visual token representations. This feature is friendly for the multi-modal dialogue task, where users may send photos of various sizes.
Given input image $\bm{I}$, Q-Former generates a fixed number of visual embeddings $\bm{H}^{\text{v}}$ (32 tokens) based on a set number of learnable query embeddings $\bm{Q}$ (32 tokens) and visual features $\widetilde{\bm{H}^{\text{v}}}$ encoded by pretrained ViT encoder (Dosovitskiy et al., 2021):
(3) $\widetilde{\bm{H}^{\text{v}}}$ $=\text{ViT}(\bm{I})$
[…]
After that, visual embeddings $\bm{H}^{\text{v}}$ are fed to the large language model as dialogue context. Since the modality representation conversion occurs entirely in the dense vector space, the image perceptron is an end-to-end model.
2.4. Image Generator
Photo-sharing multi-modal dialogue requires that the multi-modal chatbot can actively share a photo at the right moment, so we design the image generator to be responsible for the image generation function. Diffusion-based models have recently achieved impressive results in image generation tasks (Ramesh et al., 2022; Ho et al., 2020; Rombach et al., 2022). We adapt stable diffusion model (Rombach et al., 2022) in our end-to-end model, as it conducts the diffusion process in latent space not pixel space, offering higher computational efficiency and lower resource requirements. Concretely, stable diffusion consists of a clip text encoder (Radford et al., 2021) and a U-Net model (Ronneberger et al., 2015). The U-Net model iteratively performs the denoising process on the sampled gaussian noise to generate images, guided by the text representations of image text descriptions from the clip text encoder. Similar to the handling of the visual modality in the input end, the pipeline model still drives the vision model through the discrete image text description for image generation in the output end. Our end-to-end model also employs image text description to bridge the large language model and the stable diffusion model, but we propose the Straight-Through Gumbel-Softmax (Bengio et al., 2013; Jang et al., 2017; Maddison et al., 2017) and dynamic vocabulary transformation matrix to implement end-to-end gradient propagation from the image generator to the text dialogue generator.
In this section, we introduce an end-to-end image generator, which ensures gradient flow from the large language model to the stable diffusion model. First, Straight-Through Gumbel-Softmax is presented in section § 2.4.1. Then, we propose a novel dynamic vocabulary transformation matrix in section § 2.4.2, which addresses the sub-word encoding issue of vocabulary in the naive vocabulary transformation matrix. Finally, section § 2.4.3 introduces the stable diffusion model to generate images.
2.4.1. Straight-Through Gumbel-Softmax
The pipeline model completely isolates the text dialogue generation model and the image generation model and only bridges them with the discrete image text caption. For one thing, the textual signal is discrete, interrupting gradient propagation and preventing interaction between text and image generation models. For another thing, the separate training causes the image generation model to be sensitive to the text dialogue generation model. The error accumulation restricts the final performance and user experience of the multi-modal dialogue agent. Based on the straight-through gumbel-softmax (Bengio et al., 2013; Jang et al., 2017; Maddison et al., 2017), we propose an end-to-end photo-sharing multi-modal dialogue model to connect the large language model and stable diffusion model in continuous vector space not discrete textual space. Especially, our end-to-end model contributes to the image generation model not only being robust to the text dialogue model but also using powerful dialogue representations from the large language model.
Concretely, the textual dialogue model applies argmax operation on the vocabulary probability $\bm{p}$ for the next token prediction to obtain the image text caption auto-regressively, which is then delivered to the stable diffusion model to guide photo generation. The argmax and textual input are the sources of weak robustness and the isolation between text dialogue generation and image generation. Compared with the argmax operation that generates deterministic results, gumbel-softmax introduces randomness into the prediction, resulting in a sampled vector. More importantly, gumbel-softmax adapts a reparameterization trick to make the sampling process differentiable.
In particular, we feed the dialogue context $U$ and the partial response sequence of the first $i-1$ time steps $r_{<i}$ into the text dialogue generator, obtaining the token probability distribution vector $\bm{p}_{i}$ of the large language model as shown in Equation 1. Then, we sample noise vector $\bm{g}$ from standard gumbel distribution. Following (Jang et al., 2017; Maddison et al., 2017), the standard gumbel distribution is generated using inverse transform sampling on the random variable $\bm{u}$ from standard uniform distribution:
(4) $\bm{g}$ $\sim Gumbel(0,1)$
$\bm{g}$ $=-\text{log}(-\text{log}(\bm{u}))$
$\bm{u}$ $\sim Uniform(0,1)$
where $Gumbel(0,1)$ and $Uniform(0,1)$ denote standard gumbel distribution and standard uniform distribution, respectively. After that, we add the random noise into the vocabulary probability distribution of the $i$-th time step $\bm{p}_{i}$, and gain a new vocabulary probability distribution $\bm{p}_{i}^{\text{GS}}$ by softmax:
(5) $\bm{p}_{i,j}^{\text{GS}}=\frac{\text{exp}((\text{log}(\bm{p}_{i,j})+\bm{g}_{i,j})/\tau)}{\sum_{k=0}^{|V^{\text{LLM}}|}\text{exp}(\text{log}(\bm{p}_{i,k})+\bm{g}_{i,k})/\tau}$
where $\bm{p}_{i,j}$, $\bm{p}_{i,j}^{\text{GS}}$ and $\bm{g}_{i,j}$ are the $j$-th values of the vectors $\bm{p}_{i}$, $\bm{p}_{i}^{\text{GS}}$ and $\bm{g}_{i}$, respectively. $|V^{\text{LLM}}|$ indicates the size of the large language model vocabulary $V^{\text{LLM}}$. As the softmax temperature $\tau$ approaches 0, samples from the gumbel-softmax distribution become one-hot. Besides, gumbel-softmax conducts a logarithm operation on token probability $\bm{p}_{i}$, which is the result of an exponent arithmetic on similarity vector $\bm{o}_{i}$. Hence, we substitute $\text{log}(\bm{p}_{i})$ with $\bm{o}_{i}$ for computation convenience as follows:
(6) $\bm{p}_{i,j}^{\text{GS}}=\frac{\text{exp}((\bm{o}_{i,j}+\bm{g}_{i,j})/\tau)}{\sum_{k=0}^{|V^{\text{LLM}}|}\text{exp}(\bm{o}_{i,k}+\bm{g}_{i,k})/\tau}$
which is almost equivalent to Equation 5 in math and does not influence model performance in our experiments.
In order to bridge the text dialogue generator and image generator, we still need a discretized one-hot vector to indicate the predicted token in the vocabulary. Consequently, we apply hard gumbel-softmax to avoid the soft categorical probability, which achieves the exact one-hot output $\bm{r}_{i}$ by argmax and retains gradient by straight-through gradient estimator (Bengio et al., 2013) trick:
(7) $\widetilde{\bm{r}_{i}}$ $=\text{argmax}(\bm{p}_{i}^{\text{GS}})$
$\bm{r}_{i}$ $=\widetilde{\bm{r}_{i}}-\text{sg}[\bm{p}_{i}^{\text{GS}}]+\bm{p}_{i}^{\text{GS}}$
where $\text{sg}[\cdot]$ denotes the stop-gradient operation. Straight-through gradient estimator not only ensures exactly discrete vector outputs but also copies gradient from $\bm{r}_{i}$ to $\bm{p}_{i}^{\text{GS}}$. Particularly, since the added noise follows a Gumbel distribution, the distribution of the predicted token from argmax is essentially consistent with the original distribution. Finally, we aggregate one-hot token predictions from the large language model into an image text caption sequence $\bm{R}^{\text{LLM}}=(\bm{r}_{1},\cdots,\bm{r}_{m})$.
2.4.2. Dynamic Vocabulary Transformation Matrix
Vocabulary Transformation Matrix
Aiming to achieve the end-to-end photo-sharing multi-modal dialogue model, we can not convert the sampled one-hot output $\bm{R}^{\text{LLM}}=(\bm{r}_{1},\cdots,\bm{r}_{m})$ into textual caption by large language model tokenizer and vocabulary, otherwise the gradient propagation will be interrupted. Therefore we need to feed the vectorized output into the stable diffusion model directly, which will not lose differentiability. However, the vocabularies of the large language model $V^{\text{LLM}}$ and the stable diffusion model $V^{\text{SD}}$ are mismatched. To align the mismatched vocabularies, we propose vocabulary transformation matrix $\bm{M}$ to implement the conversion of one-hot character representations, which is defined as follows:
(8) $\bm{M}_{i,j}=\begin{cases}1,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{if}~{}~{}V^{\text{LLM}}[i]=V^{\text{SD}}[j],\\ 0,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{otherwise,}\end{cases}$
where $V^{\text{LLM}}[i]$ and $V^{\text{SD}}[j]$ indicate the $i$-th and $j$-th token in the vocabulary of large language model and stable diffusion model, respectively. Based on the vocabulary transformation matrix $\bm{M}$, we can acquire one-hot image text caption input with the format of one-hot vectorized representation $\bm{R}^{\text{SD}}$ for the image generation model:
(9) $\bm{R}^{\text{SD}}$ $=(\bm{r}_{1}^{\text{SD}},\cdots,\bm{r}_{n}^{\text{SD}})$
$=(\bm{r}_{1}^{\text{LLM}}\bm{M},\cdots,\bm{r}_{n}^{\text{LLM}}\bm{M})$
$=\bm{R}^{\text{LLM}}\bm{M}$
In other words, we formalize the mapping of one-hot image text caption representation from $V^{\text{LLM}}$ to $V^{\text{SD}}$ as matrix multiplication. Besides, the operation can also ensure gradient propagation between different vocabularies.
Dynamic Vocabulary Transformation Matrix
The above vocabulary transformation method is under a strong assumption: the corresponding relation of tokens in the two vocabularies is one-to-one. However, the real situation is difficult to satisfy such hypothesis, because current vocabularies usually use word piece or sentence piece tokenization. These tokenization methods split words into different tokens according to hyperparameters, corpora, tokenization algorithms, etc. Therefore, the differences between tokens from different vocabularies are large. Many tokens appear in only one vocabulary (i.e., one-to-zero), resulting in a very sparse matrix $\bm{M}$. A direct method for this issue is to apply word-level corresponding relations to construct the matrix. Concretely, we assign the mapping relation between each token in the token set of a word and all tokens in the token set of the same word, which is many-to-many relation in fact. However, if we construct the vocabulary transformation matrix on all English words before model training, $\bm{M}$ will be very dense and raise over-mapping and mis-mapping problems. In particular, most different tokens from the two vocabularies have mapping relations, which could not maintain exact and delicate token alignments.
To address this issue, we propose the dynamic vocabulary transformation matrix. We dynamically construct the vocabulary matrix $\bm{M}$ during each batch computing based on the word in the sentence, not before model training based on all English words. Constructing the vocabulary transformation matrix on only one word reduces its density and precisely ensures the mapping relationships of all tokens in the word. However, this operation is difficult for training, and computationally unfeasible, because the vocabulary transformation matrix initialization is performed repeatedly, wasting GPU computational time and memory. Therefore, we design the sentence-level dynamic vocabulary transformation matrix to balance dialogue performance and computational efficiency. Specifically, given the token set of an image text caption on the large language model vocabulary and stable diffusion model vocabulary, $\mathcal{T}^{\text{LLM}}$ and $\mathcal{T}^{\text{SD}}$, we assign token alignments to all tokens of the two sets in the dynamic vocabulary transformation matrix:
(10) $\bm{M}_{i,j}=\begin{cases}1,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{if}~{}~{}V^{\text{LLM}}[i]\in\mathcal{T}^{\text{LLM}}\bigwedge V^{\text{SD}}[j]\in\mathcal{T}^{\text{SD}},\\ 0,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{otherwise,}\end{cases}$
During the training phase, we initialize the vocabulary transformation matrix for each image caption in the batch dialogue. It is important to note that the vocabulary transformation matrix can lead to significant GPU memory overhead and limited batch size. Suppose the sizes of the large language model and stable diffusion model vocabularies are both 40,000, each matrix will occupy about $40,000*40,000*2=1,600,000,000*2=3.2G$ bytes GPU space. However, the transformation matrix is a sparse matrix factually. We hence implement it through the special sparse tensor class in numpy (Harris et al., 2020) and pytorch (Ansel et al., 2024) [^1] https://pytorch.org/docs/stable/sparse.html, which constrains GPU memory occupation to a magnitude of kilobytes (KB).
We acquire vector representation of the image text caption on the stable diffusion model vocabulary according to the proposed sentence-level dynamic vocabulary transformation matrix and Equation 9. However, the vector representation is not a standard one-hot vector format. All positions of image caption tokens in each vector are 1, not only one position. In addition, another intractable problem is that the length of the converted image caption is not adaptive to the tokenization of the stable diffusion model. To address these issues, we first use the textual image caption to get a valid one-hot vector representation $\widetilde{\bm{R}^{\text{SD}}}$ with the correct length by the stable diffusion model tokenizer. Second, we feed the converted representation $\bm{R}^{\text{SD}}$ into average pooling to eliminate sequence length dimension. Third, straight-through gradient estimator (Bengio et al., 2013) is adopted to propagate gradients again:
(11) $\bm{R}^{\text{SD}}$ $=\widetilde{\bm{R}^{\text{SD}}}-\text{sg}[\text{avg}(\bm{R}^{\text{LLM}}\bm{M})]+\text{avg}(\bm{R}^{\text{LLM}}\bm{M})$
where avg indicates average pooling. For inference, we directly feed $\widetilde{\bm{R}^{\text{SD}}}$ to the stable diffusion model because gradients is not necessary.
2.4.3. Stable Diffusion Model
The stable diffusion model (Rombach et al., 2022) is employed in our end-to-end photo-sharing multi-modal dialogue system for image generation. Concretely, the stable diffusion model consists of a clip text encoder (Radford et al., 2021) and an image U-Net model (Ronneberger et al., 2015). The former encodes the representation of image text captions, which is feed to the U-Net model. The latter iteratively performs the denoising process on the sampled gaussian noise $\bm{a}$ to generate images, guided by the text representations of image text descriptions from the clip text encoder. Finally, we exploit mean squared error (MSE) loss between the predicted image $\widehat{\bm{I}}$ and the gold standard image $\bm{I}$ to train the model:
(12) $\widehat{\bm{I}}$ $=\text{StableDiffusion}(\bm{R}^{\text{SD}},\bm{a})$
$\mathcal{L}^{\text{v}}$ $=\|\widehat{\bm{I}}-\bm{I}\|$
[…]
2.5. End-to-End Training
For the input end, Q-Former connects text and vision modality, which gives the large language model the ability to see. For the output end, the dynamic vocabulary transformation matrix integrates Straight-Through Gumbel-Softmax to implement end-to-end gradient propagation, enabling the large language model to draw images. Our end-to-end photo-sharing multi-modal dialogue realizes fluent gradient flow throughout the entire dialogue system, which can be jointly trained:
(13) $\mathcal{L}$ $=\mathcal{L}^{\text{t}}+\alpha\mathcal{L}^{\text{v}}$
where $\alpha$ is the weighted coefficient.
For the inference stage, the image perception encodes all the photos and then feeds visual representations to the text dialogue generator as dialogue contexts. The large language model generates the text response, where contents between image start token [IMG] and image end token [/IMG] are directly inputted to the stable diffusion model for quick image generation.
3. Experiments
We conduct experiments to verify the effectiveness of our proposed end-to-end model for photo-sharing multi-modal dialogue. In this section, we start with an introduction of experimental settings, then show main experiment results to verify the effectiveness of our end-to-end multi-modal dialogue model and end with some fine-grained experimental analyses to gain a deeper understanding of our approach.
3.1. Experimental Setup
3.1.1. Datasets
We conduct experiments on two datasets: PhotoChat (Zang et al., 2021) [^2] https://github.com/google-research/google-research/tree/master/multimodalchat/ and DialogCC (Lee et al., 2024) [^3] https://github.com/passing2961/DialogCC. We obtain text dialogue contents from corresponding GitHub links. For shared photos in the datasets, we download them according to URLs of images. However, some image are not accessible now due to expired URLs. We leverage stable diffusion 3 (Esser et al., 2024) [^4] https://huggingface.co/stabilityai/stable-diffusion-3-medium-diffusers to generate the missing images based on the image text captions in datasets. We employ standard train, dev and test dataset splits for model training, evaluation and testing, respectively. Table 1 lists dataset statistics. We report nine statistics from left to right: “the name of datasets”, “splits of datasets”, “the number of dialogues”, “the number of utterances”, “the number of utterances when counting consecutive utterances from the same speaker as one utterance”, “average consecutive utterances per dialogue”, “the number of images”, “the number of distinct images” and “average images per dialogue”. In particular, utterances in which photos are shared are excluded in the calculation.
Table 1. Dataset statistics. From left to right beginning from the third column, the name of each column means “the number of dialogues”, “the number of utterances”, “the number of utterances when counting consecutive utterances from the same speaker as one utterance”, “average consecutive utterances per dialogue”, “the number of images”, “the number of distinct images” and “average images per dialogue”. Utterances in which photos are shared are excluded in the calculation.
Dataset Split # Dial # Utter # CUtter U/D # Img # DImg I/D
PhotoChat (Zang et al., 2021) train 10,286 130,459 92,479 8.99 10,286 8,889 1.0
dev 1,000 12,695 9,045 9.05 1,000 1,000 1.0
test 1,000 12,841 9,089 9.09 1,000 1,000 1.0
total 12,286 155,995 110,613 9.00 12,286 10,889 1.0
DialogCC (Lee et al., 2024) train 68,402 553,990 329,969 4.82 106,198 38,743 1.55
dev 7,644 63,140 37,181 4.86 11,669 6,413 1.53
test 7,324 60,260 35,633 4.87 11,157 6,279 1.52
total 83,370 677,390 402,783 4.83 129,024 51,435 1.55
3.1.2. Hyperparameters
For the text dialogue generator, we exploit Llama-2-7b-chat-hf (Touvron et al., 2023) as the dialogue backbone. For the image perceptron, the ViT encoder and Q-Former are directly derived from the pre-trained BLIP-2 (blip2-opt-2.7b-coco) (Li et al., 2023). For the image generator, we fine-tune the stable-diffusion-2-1-base (Rombach et al., 2022), which consists of the pre-trained CLIP encoder (Radford et al., 2021) and U-Net model (Ronneberger et al., 2015). Besides, for hyper-parameters of model training, we use the AdamW algorithm with learning rate 5e-5, batch size 32, weight decay 0.01 and linear learning rate warmup over the first 1,000 steps to optimize parameters. The temperature parameter $\tau$ is gradually annealed from 1 to 0.01 in the first three epochs and remains constant in the subsequent rounds. Weight factor $\alpha$ is 1. We stop early training when the BLEU-2 and FID scores do not increase on the development set for 4 epochs. We run all experiments on Nvidia V100.
3.1.3. Evaluation
We evaluate our end-to-end model on both text and image generation for the photo-sharing multi-modal dialogue task. For text generation, we leverage BLEU-1, BLEU-2 (Papineni et al., 2002) and Rouge-L (Lin, 2004) as evaluation metrics. For image generation, we follow Sun et al. (2022) to use Inception Score (IS) (Salimans et al., 2016) and Frechet Inception Distance (FID) (Heusel et al., 2017), which are computed by torch-fidelity  [^5] https://github.com/toshas/torch-fidelity toolkit. We conduct the experiments on three different random seeds and report the average results.
3.1.4. Baselines
To verify the effectiveness of our end-to-end dialogue model, we compare results with two related baseline models:
  • Divter: Sun et al. (2022) use DialoGPT (Zhang et al., 2020) and DALL-E (Ramesh et al., 2021) to implement the pipeline-based model, where discrete image text captions bridge the text and image generation models.
  • pipeline: We implement the pipeline model for a fair comparison with our end-to-end model. This model adapts the same Llama-2 and stable-diffusion-2-1 models as the end-to-end model, but trains them separately.
3.2. Main Results
Table 2. Main results for text and image response generation. indicates the model is trained on extra resources of image generation.
Model Text Response Generation Image Response Generation
BLEU-1 BLEU-2 Rouge-L IS $\uparrow$ FID $\downarrow$
PhotoChat
Divter (Sun et al., 2022) 6.52 1.66 5.69 15.8 $\pm$ 0.6 29.16
Divter (Sun et al., 2022) 6.28 1.51 5.40  4.9 $\pm$ 0.7 262.09
pipeline 10.65 2.03 9.76  13.67 $\pm$ 1.31 79.74
end-to-end 12.08 2.94 11.00  14.47 $\pm$ 0.94 75.88
DialogCC
Divter (Sun et al., 2022) 8.08 1.89 6.71  9.19 $\pm$ 2.01 108.50
pipeline 14.18 2.93 12.21  22.99 $\pm$ 1.92 61.66
end-to-end 15.84 3.97 13.70 25.56 $\pm$ 1.79 59.63
Table 2 reports the main results of different models, including BLEU-1, BLEU-2 and Rouge-L for text response generation and IS and FID for image response generation.
First, we examine the performances of the pipeline and end-to-end model. For the PhotoChat dataset, our end-to-end model can bring significant improvements for photo-sharing multi-modal dialogue, where gains are $12.08-10.65=1.43$, $2.94-2.03=0.91$ and $11.00-9.76=1.24$ on BLUE-1, BLUE-2 and Rouge-L for text response generation, and $14.47-13.67=0.80$ and $79.74-75.88=3.86$ on IS and FID for image response generation. For the DialogCC dataset, the end-to-end model also advances the text response generation scores by $15.84-14.18=1.66$, $3.97-2.93=1.04$ and $13.70-12.21=1.49$ on BLUE-1, BLUE-2 and Rouge-L, and the image response generation scores by $25.56-22.99=2.57$ and $61.66-59.63=2.03$ on IS and FID, respectively. All enhancements verify the effectiveness of our proposed end-to-end photo-sharing multi-modal dialogue model compared with the basic pipeline model.
Second, we make comparisons between our end-to-end model and Divter, which is also a representative pipeline model. For the DialogCC dataset, our proposed end-to-end model significantly outperforms Divter, which might be due to the strong capabilities of text generation by Llama-2-7b-chat-hf and image generation by stable-diffusion-2-1-base. Concretely, the end-to-end model boosts the performances of text and image response generation significantly, leading to increases on BLUE-1 by $15.84-8.08=7.76$, BLUE-2 by $3.97-1.89=2.08$, Rouge-L by $13.70-6.71=6.99$, IS by $25.56-9.19=16.37$ and FID by $108.50-59.63=48.87$, respectively. For the PhotoChat dataset, our end-to-end model still achieves better performances compared with naive Divter, leading improvements of text response generation on BLEU-1 by $12.08-6.28=5.80$, BLEU-2 by $2.94-1.51=1.43$ and Rouge-L by $11.00-5.40=5.60$, and image response Generation on IS by $14.47-4.9=9.57$ and FID by $262.09-75.88=186.21$, respectively. Our end-to-end model reaches state-of-the-art performance compared with pipeline models, which verifies the effectiveness of our proposed dialogue system.
Third, we look at the image response generation scores. The best results of IS and FID are $15.8$ and $29.16$, which is from Divter. An important reason could be that Sun et al. (2022) employ extra image caption resources to train the model, including YFCC100M (Thomee et al., 2016) and ImageNet (Deng et al., 2009). Sampling data from the two datasets might significantly influence the scores of IS and FID, because they are computed by the neural network trained on datasets with similar data distribution.
4. Conclusion
In this paper, we proposed the first end-to-end photo-sharing multi-modal dialogue model, which consists of three modules: text dialogue generator, image perceptron and image generator. We exploited a large language model as the backbone of the text dialogue generator. For the image perceptron, Q-Former equipped the large language model with the ability to see images in the input end. For the image generator, we applied the stable diffusion model to enable the large language model to draw images in the output end. Particularly, we proposed the dynamic vocabulary transformation matrix and the straight-through gumbel-softmax technique to align the large language model and stable diffusion model and achieve end-to-end gradient propagation. Finally, we conducted experiments on PhotoChat and DialogCC datasets to verify the effectiveness of our proposed end-to-end model. Compared with two pipeline models, the end-to-end model gained state-of-the-art performances for photo-sharing multi-modal dialogue generation.
Acknowledgements.
This work was supported by the stability support plan for Shenzhen colleges and universities.

### v2

@@
 An End-to-End Model for Photo-Sharing Multi-Modal Dialogue Generation
-Abstract.
+Abstract
 Photo-Sharing Multi-modal dialogue generation requires a dialogue agent not only to generate text responses but also to share photos at the proper moment.
 Using image text caption as the bridge, a pipeline model integrates an image caption model, a text generation model, and an image generation model to handle this complex multi-modal task.
-However, representing the images with text captions may loss important visual details and information and cause error propagation in the complex dialogue system.
+However, representing the images with text captions may lose important visual details and information and cause error propagation in the complex dialogue system.
 Besides, the pipeline model isolates the three models separately because discrete image text captions hinder end-to-end gradient propagation.
 We propose the first end-to-end model for photo-sharing multi-modal dialogue generation, which integrates an image perceptron and an image generator with a large language model.
-The large language model employs the Q-Former to perceive visual images in the input end.
+The large language model employs the vision encoder to perceive visual images in the input end.
 For image generation in the output end, we propose a dynamic vocabulary transformation matrix and use straight-through and gumbel-softmax techniques to align the large language model and stable diffusion model and achieve end-to-end gradient propagation.
@@
 Compared with pipeline models, the end-to-end model gains state-of-the-art performances on various metrics of text and image generation.
-More analysis experiments also verify the effectiveness of the end-to-end model for photo-sharing multi-modal dialogue generation.
+Index Terms:
 Multi-modal dialogue, Photo-Sharing, Large language model, Stable diffusion, End-to-end model
-copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TOMMccs: Computing methodologies Natural language generationccs: Computing methodologies Discourse, dialogue and pragmaticsccs: Computing methodologies Image representationsccs: Computing methodologies Reconstruction
-1. Introduction
-Human-machine dialogue system is a crucial research and application area for the artificial intelligence community (Zhang et al., 2020; Roller et al., 2021; Adiwardana et al., 2020; Yi et al., 2024).
-With the development of vision-language representation learning, multi-modal dialogue has attracted widespread attention (Das et al., 2017; Mostafazadeh et al., 2017; Shuster et al., 2020; Zheng et al., 2022; Shuster et al., 2021).
-The multi-modal settings in these literatures are limited to adding visual modality to the input, while the output of the dialogue model remains pure textual modality.
+I Introduction
+Human-machine dialogue system is a crucial research and application area for the artificial intelligence community.
+With the development of vision-language representation learning, multi-modal human-machine dialogue has attracted widespread attention [1, 2, 3].
+However, the multi-modal settings in these literatures and recent multi-modal large language models (MLLMs) (e.g., Llama-Vision [4], Qwen-VL [5]) are limited to adding visual modality to the input, while the output of the dialogue model remains pure textual modality.
 Similar to the conversation manner of humans on instant messaging softwares (e.g., Facebook, WhatsApp, and WeChat), an engaging multi-modal dialogue agent should not only perceive and comprehend visual content but also generate images.
-Therefore, researchers have started to focus on photo-sharing multi-modal dialogue in recent years (Lee et al., 2021; Zang et al., 2021; Sun et al., 2022; Feng et al., 2023; Lee et al., 2024; Aboutalebi et al., 2024), which requires the active behavior of sharing photos in multi-turn dialogue for multi-modal chatbots.
-In order to build the photo-sharing multi-modal dialogue system, Sun et al.
-(2022) propose a pipeline model, Divter, which composes three modules: image caption model, text dialogue model (DialogGPT (Zhang et al., 2020)), and image generation model (DALL-E (Ramesh et al., 2021)).
+Therefore, researchers have started to focus on photo-sharing multi-modal dialogue in recent years [6, 7, 8, 9, 10, 11], which requires the active behavior of sharing photos in multi-turn dialogue like Figure 1 for multi-modal chatbots.
+Figure 1: Photo-sharing multi-modal dialogue.
+In order to build the photo-sharing multi-modal dialogue system, [8] proposes a pipeline model, Divter, which composes three modules: image caption model, text dialogue model, and image generation model.
 The pipeline model utilizes the image text description as the bridge to connect visual and textual modality.
-Specifically, for the input end, the text dialogue model perceives photos based on the image descriptions from the image caption model.
-For the output end, the text dialogue model generates text responses and image descriptions, which are converted into photos by the image generation model.
 Although the pipeline model can achieve photo-sharing multi-modal dialogue, the system still faces many challenges.
@@
 Any small mistake in the image descriptions will be propagated and enlarged into the generated text dialogue responses and images.
-Third, although the pipeline model includes three neural network models, they are independent and cannot achieve end-to-end gradient propagation and training.
-So the separate training of these models makes it difficult to capture the internal relationships between different modules, thereby hindering a unified end-to-end photo-sharing multi-modal dialogue model.
+Third, the pipeline model includes three neural network models, which are independent and cannot achieve end-to-end gradient propagation and training.
+So the separate training of these models makes it difficult to capture the internal relationships between different modules.
 Besides, the cascaded system has redundant functions and parameters, leading to latency and inefficiency in practical deployment.
-Recent progress in multi-modal representation alignment (Radford et al., 2021; Zhang et al., 2024; Bordes et al., 2024), large language models (LLMs) (Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023) and diffusion models (Ho et al., 2020; Rombach et al., 2022) offer potential solutions and directions for addressing these issues.
-For one thing, large language models and diffusion models have achieved remarkable results in both text and image generation tasks.
-Advanced foundation models can improve the final performance of the photo-sharing multi-modal dialogue system.
-For another thing, current multi-modal large language models integrate visual encoders with large language models, unifying the representation space of vision and text.
-This enables text generation models to directly perceive the entire image, rather than relying on image captions that may lose visual details.
-Furthermore, a single multi-modal model performing both image perception and text generation fully liberates the image caption model.
-In this work, we propose an end-to-end photo-sharing multi-modal dialogue model to alleviate the challenges in the pipeline model, resorting to recent multi-modal generation models and end-to-end gradient connection techniques (Straight-Through Estimator (Bengio et al., 2013) and Gumbel-Softmax (Jang et al., 2017; Maddison et al., 2017)).
-First, we use the large language model (Llama 2 (Touvron et al., 2023)) as the backbone of text dialogue generation.
-Second, Q-Former in BLIP-2 (Li et al., 2023) connects the vision representation encoder (ViT (Dosovitskiy et al., 2021)) with the large language model, which enables the realization of the end-to-end multi-modal model at the input end.
-Third, the stable diffusion (Rombach et al., 2022) is engaged in image generation in our photo-sharing dialogue system.
-Especially, we design a dynamic vocabulary transformation matrix to convert LLM vocab into diffusion model vocab, which integrates Straight-Through Estimator (Bengio et al., 2013) and Gumbel-Softmax (Jang et al., 2017; Maddison et al., 2017) to connect the diffusion model with the large language model and realize the end-to-end gradient connection at the output end.
-Overall, the vision encoder and the diffusion model equip the large language model with the ability to see and draw images at the same time.
-Q-Former, dynamic vocabulary transformation matrix and Straight-Through Gumbel-Softmax implement end-to-end gradient propagation and joint training.
-We conduct experiments on PhotoChat (Zang et al., 2021) and DialogCC (Lee et al., 2024) datasets to verify the effectiveness of our end-to-end photo-sharing multi-modal dialogue model.
-We employ many diverse textual and visual metrics to evaluate the performances of text and image generation, including BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), Inception Score (IS) (Salimans et al., 2016) and Frechet Inception Distance (FID) (Heusel et al., 2017).
-Our end-to-end model acquires improvements on all metrics compared with the conventional pipeline model (Divter) and LLM-based pipeline model.
-Figure 1. Photo-sharing Multi-modal dialogue.
-In conclusion, the main contributions can be summarized as follows:
-  • We propose the first end-to-end photo-sharing multi-modal dialogue model, which jointly trains the image encoder, large language model and diffusion model to acquire the abilities of image perceiving, text generation and image generation.
-  • We design a novel dynamic vocabulary transformation matrix to convert image text description from LLM to diffusion model while retaining gradients.
-  • Experimental results on diverse metrics verify the effectiveness of our end-to-end model compared with pipeline models.
-2. Model
-Figure 2.
-The overview of our end-to-end photo-sharing multi-modal dialogue generation model, which integrates an image perceptron and an image generator with a large language model.
-The large language model employs the Q-Former to perceive visual images in the input end.
+In this work, we propose an end-to-end photo-sharing multi-modal dialogue model to alleviate the challenges in the pipeline model, resorting to recent MLLMs, diffusion models [12] and end-to-end gradient connection techniques (Straight-Through Estimator [13] and Gumbel-Softmax [14, 15]).
+For one thing, we use the MLLM (i.e., Llama-Vision) as the backbone of dialogue generation, which enables the realization of the end-to-end multi-modal model at the input end.
+For another thing, the stable diffusion [12] is engaged in image generation.
+Especially, inspired by [16] and [17], we design a dynamic vocabulary transformation matrix to convert LLM vocabulary into diffusion model vocabulary, which integrates Straight-Through Estimator [13] and Gumbel-Softmax [14, 15] to connect the diffusion model with the MLLM and realize the end-to-end gradient connection at the output end.
+Overall, the MLLM and the diffusion model equip the large language model with the ability to see and draw images at the same time.
+In conclusion, the main contributions can be summarized as follows: i) We propose the first end-to-end photo-sharing multi-modal dialogue model, which jointly trains the MLLM and diffusion model to acquire the abilities of image perceiving, text generation and image generation.
+ii) We design a novel dynamic vocabulary transformation matrix and apply straight-through and gumbel-softmax techniques to convert image text description from LLM to diffusion model while retaining gradients.
+iii) Experimental results on diverse metrics verify the effectiveness of our end-to-end model compared with pipeline models  [^1] We public our code at https://github.com/guopeiming/E2E_PSDG..
+II Related Work
+With the development of vision-language representation learning in this decade, multi-modal dialogue has attracted widespread attention [1, 2, 3].
+However, conventional multi-modal dialogue tasks consider vision modality only at the input end.
+Therefore, researchers have focused on photo-sharing multi-modal dialogue in recent years [6, 7, 8, 9, 10, 11].
+[7] proposes a retrieval model to predict the image shared in the dialogue, which computes the similarity between dialogue context and image.
+[9] also models photo-sharing multi-modal dialogue by a retrieval task.
+Different from only predicting images in [7], [9] retrieves suitable textual utterances and shared photos for conversations.
+In addition to the retrieval model, [8] presents a generative model, Divter, for photo-sharing multi-modal dialogue generation.
+Specifically, Divter adapts a text dialogue model as the backbone, and perceives photos based on discrete image descriptions from an image caption model for the input end.
+For the output end, the text dialogue model generates text responses and image descriptions, which are converted into photos by an image generation model.
+Another important related work for this paper is multi-modal generation models, including image generation models and MLLMs.
+Image generation models have gained great attention in the deep learning community in this decade.
+In the beginning, researchers employ VAE and GAN models to generate image examples.
+Following, VQ-VAE [18] and VQ-GAN [19] quantize latent vectors into discrete codebook to improve VAE and GAN models.
+Recently, diffusion-based models have shown better image generation ability [20], like DALL-E-2 [21] and stable diffusion [12].
+Most MLLMs [22, 5] integrate a vision encoder into LLMs, which learn a mapping between text modality (large language model) and vision modality (vision encoder).
+However, these MLLMs acquire multi-modal capability only at the input end, employing linear projection [22], Q-Former [5] or cross-attention as [4] as the modality mapping adapter.
+A few MLLMs [23, 24] have recently implemented visual ability at both the input and output end by vision tokenizer.
+III Model
+Figure 2: The overview of our end-to-end photo-sharing multi-modal dialogue generation model, which integrates an image perceptron and an image generator with a large language model.
+The large language model employs the image perceptron to perceive visual images in the input end.
 For image generation in the output end, we propose a dynamic vocabulary transformation matrix (DVTM) and use straight-through and gumbel-softmax (STGS) techniques to align the large language model and stable diffusion model and achieve end-to-end gradient propagation.
-In this section, we start with the task formalization of photo-sharing multi-modal dialogue (§2.1), then introduce our end-to-end model, which consists of three components: text dialogue generator (§2.2), image perception (§2.3) and image generator (§2.4), end with the description of end-to-end training of our multi-modal dialogue model (§2.5).
-2.1. Task Formalization
+In this section, we start with the task formalization of photo-sharing multi-modal dialogue (§III-A), then introduce our end-to-end model shown in Figure 2, which consists of three components: text dialogue generator (§III-B), image perception (§III-C) and image generator (§III-D), end with the description of end-to-end training (§III-E).
+III-A Task Formalization
 Given a photo-sharing multi-modal dialogue dataset $\mathcal{D}=\{(U_{i},R_{i})\}_{i}^{I}$, where $\forall~{}i\in\{1,2,\dots,I\}$, $U_{i}$ is the dialogue context, $R_{i}$ is the response regarding to $U_{i}$.
-Dialogue context $U_{i}$ and response $R_{i}$ could include two modalities: textual elements (e.g., utterances) and visual elements (e.g., images).
 Concretely, for each dialogue context and response, $U_{i}=\{u_{k}\}_{k}^{K}$ and $R_{i}=\{r_{l}\}_{l}^{L}$ are sequences of multi-modal elements including textual utterances and visual images.
@@
 When given the multi-modal dialogue context $U$, the end-to-end model can generate multi-modal response $R$ including texts and images.
-2.2. Text Dialogue Generator
-In recent years, large language models have demonstrated powerful text generation ability.
-Besides, they can also acquire visual ability through extra vision modules, becoming multi-modal large language models.
-Therefore, we adapt the large language model, Llama (Touvron et al., 2023), as the backbone of our end-to-end photo-sharing multi-modal dialogue model.
-Specifically, Llama is a sequence-to-sequence generation model based on decoder-only transformer architecture, which autoregressively generates the next token according to the input sequence.
+III-B Text Dialogue Generator
 […]
-We input the dialogue context $U$ and response $R$ to the text dialogue generator as the source and target sequence for teacher-forcing training.
-First, the large language model generates the hidden state representation of the current time step $\bm{h}_{i}$, based on the dialogue context $U$ and the target sequence of the first $i-1$ time steps $r_{\textless i}$.
-Then, a linear projection layer (LPL) converts the dimension of the hidden state vector $\bm{h}_{i}$ into the size of the large language model vocabulary.
-This operation can be seen as computing the similarity vector $\bm{o}_{i}$ between $\bm{h}_{i}$ and all tokens in the vocabulary.
-Finally, we apply softmax to gain the token probability distribution $\bm{p}_{i}$ on the vocabulary of the large language model for the current time step:
-(1) $\bm{h}_{i}$ $=\text{LLM}(U,r_{\textless i})$
-$\bm{o}_{i}$ $=\text{LPL}(\bm{h}_{i})$
-$\bm{p}_{i}$ $=\text{SoftMax}(\bm{o}_{i})$
-In the end, we minimize the negative log-likelihood of the target sequence $y$ to optimize the text dialogue generator:
-(2) $\mathcal{L}^{\text{t}}=-\text{log}~{}P(R|U)=-\text{log}\prod_{i=1}^{|R|}~{}P(y_{i}|U,r_{\textless i})=-\sum_{i=1}^{|R|}\text{log}~{}\bm{p}_{i}^{r_{i}}$
-where $\mathcal{L}^{\text{t}}$ is the notation of text generation loss.
-For the inference stage, we sample the token from vocabulary probability distribution $\bm{p}$ and generate the response sequence auto-regressively.
-Concretely, the decoder-only-based model predicts the next token by encoding the dialogue context and the historical token sequence of the generated response.
-The text dialogue generator repeats the generation process until the generated response reaches the end token or the max sequence length.
-Particularly, contents between image start token [IMG] and image end token [/IMG] will be inputted to the image generator, guiding the diffusion model to generate appropriate images.
-Besides, we also feed the image text description and the corresponding vocabulary probability distribution $\bm{p}_{i}$ to the image generator during the training stage.
-Our novel gradient connection technique maintains gradient propagation between the large language model and the diffusion model, and hence builds an end-to-end model, which will represented in the following subsection 2.4.
-2.3. Image Perceptron
-For the photo in the input end, the pipeline model bridges text and vision modality by discrete image text description, which interrupts gradient computation and overlooks visual details.
-To enable the large language model to directly perceive the visual modality information of the entire image, rather than relying on image descriptions of textual modality, image perceptron attempts to connect a visual encoder to the large language model.
-There are two main approaches connecting visual encoders and large language models by inserting extra vision-language adaption layers: Q-Former (Li et al., 2023) and linear projection (Liu et al., 2023).
-We select Q-Former for our end-to-end photo-sharing multi-modal dialogue model, because it encodes images with arbitrary size to a fixed number of visual token representations.
-This feature is friendly for the multi-modal dialogue task, where users may send photos of various sizes.
-Given input image $\bm{I}$, Q-Former generates a fixed number of visual embeddings $\bm{H}^{\text{v}}$ (32 tokens) based on a set number of learnable query embeddings $\bm{Q}$ (32 tokens) and visual features $\widetilde{\bm{H}^{\text{v}}}$ encoded by pretrained ViT encoder (Dosovitskiy et al., 2021):
-(3) $\widetilde{\bm{H}^{\text{v}}}$ $=\text{ViT}(\bm{I})$
+We input the dialogue context $U$ and response $R$ to the text dialogue generator as the source and target sequence for teacher-forcing training on the vocabulary probability distribution $\bm{p}_{i}$ of the $i$-th time step:
+$\bm{p}_{i}$ $=\text{LLM}(U,r_{\textless i})$ (1)
+$\mathcal{L}^{\text{t}}$ $=-\text{log}~{}P(R|U)=-\Sigma_{i=1}^{|R|}\text{log}~{}\bm{p}_{i}^{r_{i}},$
+where $\bm{p}_{i}^{r_{i}}$ is probability of the token $r_{i}$.
+For the inference stage, LLM auto-regressively generates textual utterances or image captions.
+Particularly, the image caption and the corresponding vocabulary probability distribution $\bm{p}_{i}$ will be inputted to the image generator, which employs the diffusion model and our novel gradient connection technique to generate images and maintain gradient propagation.
+III-C Image Perceptron
 […]
-After that, visual embeddings $\bm{H}^{\text{v}}$ are fed to the large language model as dialogue context.
-Since the modality representation conversion occurs entirely in the dense vector space, the image perceptron is an end-to-end model.
-2.4. Image Generator
-Photo-sharing multi-modal dialogue requires that the multi-modal chatbot can actively share a photo at the right moment, so we design the image generator to be responsible for the image generation function.
-Diffusion-based models have recently achieved impressive results in image generation tasks (Ramesh et al., 2022; Ho et al., 2020; Rombach et al., 2022).
-We adapt stable diffusion model (Rombach et al., 2022) in our end-to-end model, as it conducts the diffusion process in latent space not pixel space, offering higher computational efficiency and lower resource requirements.
-Concretely, stable diffusion consists of a clip text encoder (Radford et al., 2021) and a U-Net model (Ronneberger et al., 2015).
-The U-Net model iteratively performs the denoising process on the sampled gaussian noise to generate images, guided by the text representations of image text descriptions from the clip text encoder.
+III-D Image Generator
+Photo-sharing multi-modal dialogue requires the multi-modal chatbot to share a photo at the right moment actively, so we design the image generator to be responsible for the image generation function.
+We adapt the stable diffusion model [12] in our end-to-end system for its impressive generation performance and computational efficiency.
 Similar to the handling of the visual modality in the input end, the pipeline model still drives the vision model through the discrete image text description for image generation in the output end.
-Our end-to-end model also employs image text description to bridge the large language model and the stable diffusion model, but we propose the Straight-Through Gumbel-Softmax (Bengio et al., 2013; Jang et al., 2017; Maddison et al., 2017) and dynamic vocabulary transformation matrix to implement end-to-end gradient propagation from the image generator to the text dialogue generator.
-In this section, we introduce an end-to-end image generator, which ensures gradient flow from the large language model to the stable diffusion model.
-First, Straight-Through Gumbel-Softmax is presented in section § 2.4.1.
-Then, we propose a novel dynamic vocabulary transformation matrix in section § 2.4.2, which addresses the sub-word encoding issue of vocabulary in the naive vocabulary transformation matrix.
-Finally, section § 2.4.3 introduces the stable diffusion model to generate images.
-2.4.1. Straight-Through Gumbel-Softmax
+Our end-to-end model also employs image text description to bridge the large language model and the stable diffusion model, but we propose the Straight-Through Gumbel-Softmax [13, 14, 15] and dynamic vocabulary transformation matrix to implement end-to-end gradient propagation from the image generator to the text dialogue generator.
+III-D1 Straight-Through Gumbel-Softmax
 The pipeline model completely isolates the text dialogue generation model and the image generation model and only bridges them with the discrete image text caption.
-For one thing, the textual signal is discrete, interrupting gradient propagation and preventing interaction between text and image generation models.
-For another thing, the separate training causes the image generation model to be sensitive to the text dialogue generation model.
-The error accumulation restricts the final performance and user experience of the multi-modal dialogue agent.
-Based on the straight-through gumbel-softmax (Bengio et al., 2013; Jang et al., 2017; Maddison et al., 2017), we propose an end-to-end photo-sharing multi-modal dialogue model to connect the large language model and stable diffusion model in continuous vector space not discrete textual space.
-Especially, our end-to-end model contributes to the image generation model not only being robust to the text dialogue model but also using powerful dialogue representations from the large language model.
-Concretely, the textual dialogue model applies argmax operation on the vocabulary probability $\bm{p}$ for the next token prediction to obtain the image text caption auto-regressively, which is then delivered to the stable diffusion model to guide photo generation.
+Based on the straight-through gumbel-softmax [13, 14, 15], we propose an end-to-end model to connect the LLM and stable diffusion model in continuous vector space not discrete textual space.
+Concretely, the text dialogue generator applies argmax operation to generate an image caption, which is then delivered to the stable diffusion model to guide photo generation.
 The argmax and textual input are the sources of weak robustness and the isolation between text dialogue generation and image generation.
-Compared with the argmax operation that generates deterministic results, gumbel-softmax introduces randomness into the prediction, resulting in a sampled vector.
+Compared with the argmax operation, gumbel-softmax introduces randomness into the prediction, resulting in a sampled vector.
 More importantly, gumbel-softmax adapts a reparameterization trick to make the sampling process differentiable.
-In particular, we feed the dialogue context $U$ and the partial response sequence of the first $i-1$ time steps $r_{<i}$ into the text dialogue generator, obtaining the token probability distribution vector $\bm{p}_{i}$ of the large language model as shown in Equation 1.
-Then, we sample noise vector $\bm{g}$ from standard gumbel distribution.
-Following (Jang et al., 2017; Maddison et al., 2017), the standard gumbel distribution is generated using inverse transform sampling on the random variable $\bm{u}$ from standard uniform distribution:
-(4) $\bm{g}$ $\sim Gumbel(0,1)$
-$\bm{g}$ $=-\text{log}(-\text{log}(\bm{u}))$
-$\bm{u}$ $\sim Uniform(0,1)$
-where $Gumbel(0,1)$ and $Uniform(0,1)$ denote standard gumbel distribution and standard uniform distribution, respectively.
-After that, we add the random noise into the vocabulary probability distribution of the $i$-th time step $\bm{p}_{i}$, and gain a new vocabulary probability distribution $\bm{p}_{i}^{\text{GS}}$ by softmax:
-(5) $\bm{p}_{i,j}^{\text{GS}}=\frac{\text{exp}((\text{log}(\bm{p}_{i,j})+\bm{g}_{i,j})/\tau)}{\sum_{k=0}^{|V^{\text{LLM}}|}\text{exp}(\text{log}(\bm{p}_{i,k})+\bm{g}_{i,k})/\tau}$
+Specifically, we first sample random variable $\bm{g}$ and add it into the vocabulary probability distribution $\bm{p}_{i}$ to gain a new one $\bm{p}_{i}^{\text{GS}}$.
+Then we apply hard gumbel-softmax to avoid the soft categorical probability, which achieves the exact one-hot output $\bm{r}_{i}$ by argmax and retains gradient by straight-through gradient estimator [13]:
+$\bm{g}$ $\sim Gumbel(0,1)$ (2)
+$\bm{p}_{i,j}^{\text{GS}}$ $=\frac{\text{exp}((\text{log}(\bm{p}_{i,j})+\bm{g}_{i,j})/\tau)}{\sum_{k=0}^{|V^{\text{LLM}}|}\text{exp}(\text{log}(\bm{p}_{i,k})+\bm{g}_{i,k})/\tau}$
+$\bm{r}_{i}$ $=\text{argmax}(\bm{p}_{i}^{\text{GS}})-\text{sg}[\bm{p}_{i}^{\text{GS}}]+\bm{p}_{i}^{\text{GS}}.$
 where $\bm{p}_{i,j}$, $\bm{p}_{i,j}^{\text{GS}}$ and $\bm{g}_{i,j}$ are the $j$-th values of the vectors $\bm{p}_{i}$, $\bm{p}_{i}^{\text{GS}}$ and $\bm{g}_{i}$, respectively.
-$|V^{\text{LLM}}|$ indicates the size of the large language model vocabulary $V^{\text{LLM}}$.
-As the softmax temperature $\tau$ approaches 0, samples from the gumbel-softmax distribution become one-hot.
-Besides, gumbel-softmax conducts a logarithm operation on token probability $\bm{p}_{i}$, which is the result of an exponent arithmetic on similarity vector $\bm{o}_{i}$.
-Hence, we substitute $\text{log}(\bm{p}_{i})$ with $\bm{o}_{i}$ for computation convenience as follows:
-(6) $\bm{p}_{i,j}^{\text{GS}}=\frac{\text{exp}((\bm{o}_{i,j}+\bm{g}_{i,j})/\tau)}{\sum_{k=0}^{|V^{\text{LLM}}|}\text{exp}(\bm{o}_{i,k}+\bm{g}_{i,k})/\tau}$
-which is almost equivalent to Equation 5 in math and does not influence model performance in our experiments.
-In order to bridge the text dialogue generator and image generator, we still need a discretized one-hot vector to indicate the predicted token in the vocabulary.
-Consequently, we apply hard gumbel-softmax to avoid the soft categorical probability, which achieves the exact one-hot output $\bm{r}_{i}$ by argmax and retains gradient by straight-through gradient estimator (Bengio et al., 2013) trick:
-(7) $\widetilde{\bm{r}_{i}}$ $=\text{argmax}(\bm{p}_{i}^{\text{GS}})$
-$\bm{r}_{i}$ $=\widetilde{\bm{r}_{i}}-\text{sg}[\bm{p}_{i}^{\text{GS}}]+\bm{p}_{i}^{\text{GS}}$
-where $\text{sg}[\cdot]$ denotes the stop-gradient operation.
-Straight-through gradient estimator not only ensures exactly discrete vector outputs but also copies gradient from $\bm{r}_{i}$ to $\bm{p}_{i}^{\text{GS}}$.
-Particularly, since the added noise follows a Gumbel distribution, the distribution of the predicted token from argmax is essentially consistent with the original distribution.
+$|V^{\text{LLM}}|$ indicates the size of the LLM vocabulary $V^{\text{LLM}}$.
+As the temperature $\tau$ approaches 0, samples from the gumbel-softmax distribution become one-hot.
+$\text{sg}[\cdot]$ is the stop-gradient operation for straight-through gradient estimator, which not only ensures exactly discrete vector outputs but also copies gradient from $\bm{r}_{i}$ to $\bm{p}_{i}^{\text{GS}}$.
 Finally, we aggregate one-hot token predictions from the large language model into an image text caption sequence $\bm{R}^{\text{LLM}}=(\bm{r}_{1},\cdots,\bm{r}_{m})$.
-2.4.2. Dynamic Vocabulary Transformation Matrix
+III-D2 Dynamic Vocabulary Transformation Matrix
 Vocabulary Transformation Matrix
-Aiming to achieve the end-to-end photo-sharing multi-modal dialogue model, we can not convert the sampled one-hot output $\bm{R}^{\text{LLM}}=(\bm{r}_{1},\cdots,\bm{r}_{m})$ into textual caption by large language model tokenizer and vocabulary, otherwise the gradient propagation will be interrupted.
-Therefore we need to feed the vectorized output into the stable diffusion model directly, which will not lose differentiability.
-However, the vocabularies of the large language model $V^{\text{LLM}}$ and the stable diffusion model $V^{\text{SD}}$ are mismatched.
+Aiming to achieve the end-to-end model, we can not convert the sampled one-hot output $\bm{R}^{\text{LLM}}$ into textual caption by LLM tokenizer and vocabulary, otherwise the gradient propagation will be interrupted.
+Therefore we need to feed the vectorized output into the stable diffusion model directly.
+However, the vocabularies of the LLM $V^{\text{LLM}}$ and the stable diffusion model $V^{\text{SD}}$ are mismatched.
 To align the mismatched vocabularies, we propose vocabulary transformation matrix $\bm{M}$ to implement the conversion of one-hot character representations, which is defined as follows:
-(8) $\bm{M}_{i,j}=\begin{cases}1,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{if}~{}~{}V^{\text{LLM}}[i]=V^{\text{SD}}[j],\\ 0,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{otherwise,}\end{cases}$
-where $V^{\text{LLM}}[i]$ and $V^{\text{SD}}[j]$ indicate the $i$-th and $j$-th token in the vocabulary of large language model and stable diffusion model, respectively.
-Based on the vocabulary transformation matrix $\bm{M}$, we can acquire one-hot image text caption input with the format of one-hot vectorized representation $\bm{R}^{\text{SD}}$ for the image generation model:
-(9) $\bm{R}^{\text{SD}}$ $=(\bm{r}_{1}^{\text{SD}},\cdots,\bm{r}_{n}^{\text{SD}})$
-$=(\bm{r}_{1}^{\text{LLM}}\bm{M},\cdots,\bm{r}_{n}^{\text{LLM}}\bm{M})$
-$=\bm{R}^{\text{LLM}}\bm{M}$
-In other words, we formalize the mapping of one-hot image text caption representation from $V^{\text{LLM}}$ to $V^{\text{SD}}$ as matrix multiplication.
-Besides, the operation can also ensure gradient propagation between different vocabularies.
+$\bm{M}_{i,j}=\begin{cases}1,~{}~{}~{}~{}~{}~{}~{}\text{if}~{}~{}V^{\text{LLM}}[i]=V^{\text{SD}}[j],\\ 0,~{}~{}~{}~{}~{}~{}~{}\text{otherwise,}\end{cases}$ (3)
+where $V^{\text{LLM}}[i]$ and $V^{\text{SD}}[j]$ indicate the $i$-th and $j$-th token in the vocabulary of LLM and stable diffusion model, respectively.
+Based on $\bm{M}$, we formalize the mapping of one-hot image text caption representation from $V^{\text{LLM}}$ to $V^{\text{SD}}$ as matrix multiplication (i.e., $\bm{R}^{\text{SD}}=\bm{R}^{\text{LLM}}\bm{M}$), which can ensure gradient propagation between different vocabularies.
 Dynamic Vocabulary Transformation Matrix
@@
 Specifically, given the token set of an image text caption on the large language model vocabulary and stable diffusion model vocabulary, $\mathcal{T}^{\text{LLM}}$ and $\mathcal{T}^{\text{SD}}$, we assign token alignments to all tokens of the two sets in the dynamic vocabulary transformation matrix:
-(10) $\bm{M}_{i,j}=\begin{cases}1,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{if}~{}~{}V^{\text{LLM}}[i]\in\mathcal{T}^{\text{LLM}}\bigwedge V^{\text{SD}}[j]\in\mathcal{T}^{\text{SD}},\\ 0,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{otherwise,}\end{cases}$
+$\bm{M}_{i,j}=\begin{cases}1,~{}~{}\text{if}~{}~{}V^{\text{LLM}}[i]\in\mathcal{T}^{\text{LLM}}\bigwedge V^{\text{SD}}[j]\in\mathcal{T}^{\text{SD}},\\ 0,~{}~{}\text{otherwise,}\end{cases}$ (4)
 During the training phase, we initialize the vocabulary transformation matrix for each image caption in the batch dialogue.
-It is important to note that the vocabulary transformation matrix can lead to significant GPU memory overhead and limited batch size.
-Suppose the sizes of the large language model and stable diffusion model vocabularies are both 40,000, each matrix will occupy about $40,000*40,000*2=1,600,000,000*2=3.2G$ bytes GPU space.
-However, the transformation matrix is a sparse matrix factually.
-We hence implement it through the special sparse tensor class in numpy (Harris et al., 2020) and pytorch (Ansel et al., 2024) [^1] https://pytorch.org/docs/stable/sparse.html, which constrains GPU memory occupation to a magnitude of kilobytes (KB).
-We acquire vector representation of the image text caption on the stable diffusion model vocabulary according to the proposed sentence-level dynamic vocabulary transformation matrix and Equation 9.
-However, the vector representation is not a standard one-hot vector format.
+It is important to note that the vocabulary transformation matrix can lead to significant GPU memory overhead.
+Suppose the sizes of the LLM and stable diffusion model vocabularies are both 40,000, each matrix will occupy about $40,000*40,000*2=1,600,000,000*2=3.2G$ bytes GPU space.
+However, the transformation matrix is sparse factually.
+We hence implement it through the special sparse tensor class in numpy and pytorch, which constrains GPU memory occupation to a magnitude of kilobytes (KB).
+Although we acquire the vector representation of the image caption, it is not a standard one-hot vector format.
 All positions of image caption tokens in each vector are 1, not only one position.
@@
 Second, we feed the converted representation $\bm{R}^{\text{SD}}$ into average pooling to eliminate sequence length dimension.
-Third, straight-through gradient estimator (Bengio et al., 2013) is adopted to propagate gradients again:
-(11) $\bm{R}^{\text{SD}}$ $=\widetilde{\bm{R}^{\text{SD}}}-\text{sg}[\text{avg}(\bm{R}^{\text{LLM}}\bm{M})]+\text{avg}(\bm{R}^{\text{LLM}}\bm{M})$
-where avg indicates average pooling. For inference, we directly feed $\widetilde{\bm{R}^{\text{SD}}}$ to the stable diffusion model because gradients is not necessary.
-2.4.3. Stable Diffusion Model
-The stable diffusion model (Rombach et al., 2022) is employed in our end-to-end photo-sharing multi-modal dialogue system for image generation.
-Concretely, the stable diffusion model consists of a clip text encoder (Radford et al., 2021) and an image U-Net model (Ronneberger et al., 2015).
-The former encodes the representation of image text captions, which is feed to the U-Net model.
-The latter iteratively performs the denoising process on the sampled gaussian noise $\bm{a}$ to generate images, guided by the text representations of image text descriptions from the clip text encoder.
-Finally, we exploit mean squared error (MSE) loss between the predicted image $\widehat{\bm{I}}$ and the gold standard image $\bm{I}$ to train the model:
-(12) $\widehat{\bm{I}}$ $=\text{StableDiffusion}(\bm{R}^{\text{SD}},\bm{a})$
-$\mathcal{L}^{\text{v}}$ $=\|\widehat{\bm{I}}-\bm{I}\|$
+Third, straight-through gradient estimator [13] is adopted to propagate gradients again:
+$\bm{R}^{\text{SD}}$ $=\widetilde{\bm{R}^{\text{SD}}}-\text{sg}[\text{avg}(\bm{R}^{\text{LLM}}\bm{M})]+\text{avg}(\bm{R}^{\text{LLM}}\bm{M})$ (5)
+where avg indicates average pooling.
+Finally, we employ the stable diffusion model [12] for image generation based on $\bm{R}^{\text{SD}}$.
+For inference, we directly feed $\widetilde{\bm{R}^{\text{SD}}}$ to the stable diffusion model because gradients is not necessary.
+III-E End-to-End Training
+For the input end, MLLM connects text and vision modality, which gives the LLM the ability to see.
+For the output end, the dynamic vocabulary transformation matrix integrates Straight-Through Gumbel-Softmax to implement end-to-end gradient propagation, enabling the LLM to draw images.
+Our end-to-end model realizes fluent gradient flow throughout the entire dialogue system, which can be jointly trained:
+$\mathcal{L}$ $=\mathcal{L}^{\text{t}}+\alpha\mathcal{L}^{\text{v}},$ (6)
+where $\alpha$ is the weighted coefficient, and $\mathcal{L}^{\text{v}}$ is the loss of the stable diffusion model for image generation.
+IV Experiments
+IV-A Experimental Setup
+IV-A1 Datasets
+We conduct experiments on two datasets: PhotoChat [7] and DialogCC [10].
+We obtain text dialogue contents from corresponding GitHub repos.
+For shared photos in the datasets, we download them according to the URLs of images.
+However, some image are not accessible now due to expired URLs.
+We leverage stable diffusion 3 [12] to generate the missing images based on the image text captions in datasets.
+We employ standard train, dev and test dataset splits for model training, evaluation and testing, respectively.
+IV-A2 Hyperparameters
+For the text dialogue generator and image perceptron in MLLM, we exploit Llama-Vision [4] as the backbone.
+For the image generator, we fine-tune the stable-diffusion-3-medium [12].
+Besides, for hyper-parameters of model training, we use the AdamW algorithm with learning rate 5e-5, batch size 32, weight decay 0.01 and linear learning rate warmup over the first 1,000 steps to optimize parameters.
+The temperature parameter $\tau$ is gradually annealed from 1 to 0.0001 in the first three epochs and remains constant in the subsequent rounds.
+Weight factor $\alpha$ is 1.
+We run all experiments on Nvidia V100.
+IV-A3 Evaluation
+We evaluate our end-to-end model on both text and image generation for the photo-sharing multi-modal dialogue task.
+For text generation, we leverage BLEU-1, BLEU-2 and Rouge-L as evaluation metrics.
+For image generation, we use Inception Score (IS) and Frechet Inception Distance (FID).
+We conduct the experiments on three different random seeds and report the average results.
+IV-A4 Baselines
+To verify the effectiveness of our end-to-end model, we compare results with two related baseline models: 1) Divter [8] uses DialoGPT and DALL-E to implement the pipeline-based model.
+2) pipeline adapts the same Llama-3 and stable-diffusion-3-medium as the end-to-end model, but trains them separately.
+3) GPT-4 [25] and 4) Chameleon [23] are two representative end-to-end MLLMs.
+IV-B Main Results
+TABLE I: Main results for text and image response generation. indicates the model is trained on extra resources of image generation.
+Model Text Generation Image Generation
+BL-1 BL-2 Ro-L IS $\uparrow$ FID $\downarrow$
+PhotoChat
+Divter [8] 6.52 1.66 5.69 15.8 $\pm$ 0.6 29.16
+Divter [8] 6.28 1.51 5.40 4.9 $\pm$ 0.7 262.09
+GPT-4 [25] 5.78 1.02 4.97 12.66 $\pm$ 1.05 131.40
+Chameleon [23] 8.62 1.73 7.88 6.04 $\pm$ 1.63 177.79
+pipeline 10.68 2.10 9.74 13.71 $\pm$ 1.35 79.94
+end-to-end 12.14 2.98 11.12 14.59 $\pm$ 0.91 75.84
+DialogCC
+Divter [8] 8.08 1.89 6.71 9.19 $\pm$ 2.01 108.50
+GPT-4 [25] 6.13 0.87 4.04 11.50 $\pm$ 1.97 81.56
+Chameleon [23] 11.84 2.07 8.32 10.08 $\pm$ 1.34 101.72
+pipeline 14.21 2.95 12.32 23.11 $\pm$ 1.85 61.38
+end-to-end 15.87 4.02 13.80 25.61 $\pm$ 1.66 59.40
+TABLE I reports the main results of different models.
+First, we examine the performances of the pipeline and end-to-end model.
+For the PhotoChat dataset, our end-to-end model can bring significant improvements for photo-sharing multi-modal dialogue, where gains are $12.14-10.68=1.46$, $2.98-2.10=0.88$ and $11.12-9.74=1.38$ on BLUE-1, BLUE-2 and Rouge-L for text response generation, and $14.59-13.71=0.88$ and $79.94-75.84=4.10$ on IS and FID for image response generation.
+For the DialogCC dataset, the end-to-end model also advances the text response generation scores by $1.66$, $1.07$ and $1.48$ on BLUE-1, BLUE-2 and Rouge-L, and the image response generation scores by $2.50$ and $1.98$ on IS and FID, respectively.
+All enhancements verify the effectiveness of our proposed end-to-end model compared with the basic pipeline model.
+Second, we make comparisons between our end-to-end model and Divter, which is also a representative pipeline model.
+For the DialogCC dataset, our end-to-end model significantly outperforms Divter, which might be due to the strong capabilities of text generation by Llama-3 and image generation by stable-diffusion-3.
+Concretely, the end-to-end model boosts the performances of text and image response generation significantly, leading to increases on BLUE-1 by $7.79$, BLUE-2 by $2.13$, Rouge-L by $7.09$, IS by $16.42$ and FID by $49.10$, respectively.
+For the PhotoChat dataset, our end-to-end model still achieves better performances compared with naive Divter.
+However, the best results of IS and FID are $15.8$ and $29.16$, which is from Divter.
+An important reason could be that [8] employs extra image caption resources to train the model.
+Sampling data from the two datasets might significantly influence the scores of IS and FID, because they are computed by the neural network trained on datasets with similar data distribution.
+Third, we observe the results of other end-to-end multi-modal models.
+The performance of GPT-4 is limited, because it is a close-source model and could not be fine-tuned on specific datasets.
+We find its text responses are usually too long, but the ground-truths are oral expressions and short.
+For Chameleon, image generation results underperform our end-to-end model.
+The reason could be that Chameleon opens only the text transformer parameter weights, but masks image generation parameter weights.
+IV-C Analyses
+IV-C1 Ablation Experiment.
+TABLE II: Ablation experiment.
+Model Text Generation Image Generation
+BL-1 BL-2 Ro-L IS $\uparrow$ FID $\downarrow$
+end-to-end 15.87 4.02 13.80 25.61 $\pm$ 1.66 59.40
+-Image Perceptron -1.36 -0.78 -1.18 -1.29 +0.70
+-Image Generator -0.37 -0.21 -0.42 -1.94 +1.47
+The core of our end-to-end model is to replace the discrete image text descriptions in the pipeline model with dense vector representations, therefore gradient can propagate throughout the whole model without interruption.
+We conduct an ablation experiment to show the influence of the proposed end-to-end techniques.
+TABLE II reports ablation experiment results.
+-Image Perceptron means the end-to-end model ablates dense visual representations from the image perceptron and feeds discrete image text captions into the text dialogue generator.
+-Image Generator means the text dialogue generator outputs discrete image text description for the diffusion model, which ablates the dynamic vocabulary transformation matrix and straight-through gumbel-softmax.
+The two ablations significantly decrease the text and image response generation results, verifying the effectiveness of our proposed end-to-end model.
+IV-C2 Response Generation Performances for Different Speakers.
+ppe2e13.2713.2713.2713.2714.7314.7314.7314.7315.415.415.415.417.0117.0117.0117.01BLUE-1ppe2e2.862.862.862.863.813.813.813.813.143.143.143.144.214.214.214.21BLUE-2ppe2e11.8911.8911.8911.8913.0713.0713.0713.0712.712.712.712.715.1215.1215.1215.12Rouge-Lpp speaker Ae2e speaker App speaker Be2e speaker Bppe2e21.9521.9521.9521.9524.2324.2324.2324.2323.6623.6623.6623.6628.1928.1928.1928.19ISppe2e65.1865.1865.1865.1863.6663.6663.6663.6657.3957.3957.3957.3955.0255.0255.0255.02FID
+Figure 3: Response generation performances for different speakers.
+There are two speakers in our photo-sharing multi-modal dialogue.
+We examine the text and image response generation performances of the pipeline (pp) and end-to-end (e2e) model for the two speakers respectively.
+Figure 3 illustrates the experimental results.
+Regardless of the pipeline model and end-to-end model, response performances of speaker B are superior to speaker A.
+The reason might be from the instruction tuning stage of large language models, where the loss function is computed on the user side.
+In our experiments, we employ the chat mode of large language models, and take the assistant side as speaker A, while the user side as speaker B.
+Additionally, Figure 3 also demonstrates that the end-to-end model advances all results compared with the pipeline model, verifying the effectiveness of our end-to-end model.
+IV-C3 Image Response Generation Performance for Different Gumbel-Softmax Temperature.
 […]
-2.5. End-to-End Training
-For the input end, Q-Former connects text and vision modality, which gives the large language model the ability to see.
-For the output end, the dynamic vocabulary transformation matrix integrates Straight-Through Gumbel-Softmax to implement end-to-end gradient propagation, enabling the large language model to draw images.
-Our end-to-end photo-sharing multi-modal dialogue realizes fluent gradient flow throughout the entire dialogue system, which can be jointly trained:
-(13) $\mathcal{L}$ $=\mathcal{L}^{\text{t}}+\alpha\mathcal{L}^{\text{v}}$
-where $\alpha$ is the weighted coefficient.
-For the inference stage, the image perception encodes all the photos and then feeds visual representations to the text dialogue generator as dialogue contexts.
-The large language model generates the text response, where contents between image start token [IMG] and image end token [/IMG] are directly inputted to the stable diffusion model for quick image generation.
-3. Experiments
-We conduct experiments to verify the effectiveness of our proposed end-to-end model for photo-sharing multi-modal dialogue.
-In this section, we start with an introduction of experimental settings, then show main experiment results to verify the effectiveness of our end-to-end multi-modal dialogue model and end with some fine-grained experimental analyses to gain a deeper understanding of our approach.
-3.1. Experimental Setup
-3.1.1. Datasets
-We conduct experiments on two datasets: PhotoChat (Zang et al., 2021) [^2] https://github.com/google-research/google-research/tree/master/multimodalchat/ and DialogCC (Lee et al., 2024) [^3] https://github.com/passing2961/DialogCC.
-We obtain text dialogue contents from corresponding GitHub links.
-For shared photos in the datasets, we download them according to URLs of images.
-However, some image are not accessible now due to expired URLs.
-We leverage stable diffusion 3 (Esser et al., 2024) [^4] https://huggingface.co/stabilityai/stable-diffusion-3-medium-diffusers to generate the missing images based on the image text captions in datasets.
-We employ standard train, dev and test dataset splits for model training, evaluation and testing, respectively.
-Table 1 lists dataset statistics.
-We report nine statistics from left to right: “the name of datasets”, “splits of datasets”, “the number of dialogues”, “the number of utterances”, “the number of utterances when counting consecutive utterances from the same speaker as one utterance”, “average consecutive utterances per dialogue”, “the number of images”, “the number of distinct images” and “average images per dialogue”.
-In particular, utterances in which photos are shared are excluded in the calculation.
-Table 1.
-Dataset statistics.
-From left to right beginning from the third column, the name of each column means “the number of dialogues”, “the number of utterances”, “the number of utterances when counting consecutive utterances from the same speaker as one utterance”, “average consecutive utterances per dialogue”, “the number of images”, “the number of distinct images” and “average images per dialogue”.
-Utterances in which photos are shared are excluded in the calculation.
-Dataset Split # Dial # Utter # CUtter U/D # Img # DImg I/D
-PhotoChat (Zang et al., 2021) train 10,286 130,459 92,479 8.99 10,286 8,889 1.0
-dev 1,000 12,695 9,045 9.05 1,000 1,000 1.0
-test 1,000 12,841 9,089 9.09 1,000 1,000 1.0
-total 12,286 155,995 110,613 9.00 12,286 10,889 1.0
-DialogCC (Lee et al., 2024) train 68,402 553,990 329,969 4.82 106,198 38,743 1.55
-dev 7,644 63,140 37,181 4.86 11,669 6,413 1.53
-test 7,324 60,260 35,633 4.87 11,157 6,279 1.52
-total 83,370 677,390 402,783 4.83 129,024 51,435 1.55
-3.1.2. Hyperparameters
-For the text dialogue generator, we exploit Llama-2-7b-chat-hf (Touvron et al., 2023) as the dialogue backbone.
-For the image perceptron, the ViT encoder and Q-Former are directly derived from the pre-trained BLIP-2 (blip2-opt-2.7b-coco) (Li et al., 2023).
-For the image generator, we fine-tune the stable-diffusion-2-1-base (Rombach et al., 2022), which consists of the pre-trained CLIP encoder (Radford et al., 2021) and U-Net model (Ronneberger et al., 2015).
-Besides, for hyper-parameters of model training, we use the AdamW algorithm with learning rate 5e-5, batch size 32, weight decay 0.01 and linear learning rate warmup over the first 1,000 steps to optimize parameters.
-The temperature parameter $\tau$ is gradually annealed from 1 to 0.01 in the first three epochs and remains constant in the subsequent rounds.
-Weight factor $\alpha$ is 1.
-We stop early training when the BLEU-2 and FID scores do not increase on the development set for 4 epochs.
-We run all experiments on Nvidia V100.
-3.1.3. Evaluation
-We evaluate our end-to-end model on both text and image generation for the photo-sharing multi-modal dialogue task.
-For text generation, we leverage BLEU-1, BLEU-2 (Papineni et al., 2002) and Rouge-L (Lin, 2004) as evaluation metrics.
-For image generation, we follow Sun et al.
-(2022) to use Inception Score (IS) (Salimans et al., 2016) and Frechet Inception Distance (FID) (Heusel et al., 2017), which are computed by torch-fidelity  [^5] https://github.com/toshas/torch-fidelity toolkit.
-We conduct the experiments on three different random seeds and report the average results.
-3.1.4. Baselines
-To verify the effectiveness of our end-to-end dialogue model, we compare results with two related baseline models:
-  • Divter: Sun et al.
-(2022) use DialoGPT (Zhang et al., 2020) and DALL-E (Ramesh et al., 2021) to implement the pipeline-based model, where discrete image text captions bridge the text and image generation models.
-  • pipeline: We implement the pipeline model for a fair comparison with our end-to-end model.
-This model adapts the same Llama-2 and stable-diffusion-2-1 models as the end-to-end model, but trains them separately.
-3.2. Main Results
-Table 2. Main results for text and image response generation. indicates the model is trained on extra resources of image generation.
-Model Text Response Generation Image Response Generation
-BLEU-1 BLEU-2 Rouge-L IS $\uparrow$ FID $\downarrow$
-PhotoChat
-Divter (Sun et al., 2022) 6.52 1.66 5.69 15.8 $\pm$ 0.6 29.16
-Divter (Sun et al., 2022) 6.28 1.51 5.40  4.9 $\pm$ 0.7 262.09
-pipeline 10.65 2.03 9.76  13.67 $\pm$ 1.31 79.74
-end-to-end 12.08 2.94 11.00  14.47 $\pm$ 0.94 75.88
-DialogCC
-Divter (Sun et al., 2022) 8.08 1.89 6.71  9.19 $\pm$ 2.01 108.50
-pipeline 14.18 2.93 12.21  22.99 $\pm$ 1.92 61.66
-end-to-end 15.84 3.97 13.70 25.56 $\pm$ 1.79 59.63
-Table 2 reports the main results of different models, including BLEU-1, BLEU-2 and Rouge-L for text response generation and IS and FID for image response generation.
-First, we examine the performances of the pipeline and end-to-end model.
-For the PhotoChat dataset, our end-to-end model can bring significant improvements for photo-sharing multi-modal dialogue, where gains are $12.08-10.65=1.43$, $2.94-2.03=0.91$ and $11.00-9.76=1.24$ on BLUE-1, BLUE-2 and Rouge-L for text response generation, and $14.47-13.67=0.80$ and $79.74-75.88=3.86$ on IS and FID for image response generation.
-For the DialogCC dataset, the end-to-end model also advances the text response generation scores by $15.84-14.18=1.66$, $3.97-2.93=1.04$ and $13.70-12.21=1.49$ on BLUE-1, BLUE-2 and Rouge-L, and the image response generation scores by $25.56-22.99=2.57$ and $61.66-59.63=2.03$ on IS and FID, respectively.
-All enhancements verify the effectiveness of our proposed end-to-end photo-sharing multi-modal dialogue model compared with the basic pipeline model.
-Second, we make comparisons between our end-to-end model and Divter, which is also a representative pipeline model.
-For the DialogCC dataset, our proposed end-to-end model significantly outperforms Divter, which might be due to the strong capabilities of text generation by Llama-2-7b-chat-hf and image generation by stable-diffusion-2-1-base.
-Concretely, the end-to-end model boosts the performances of text and image response generation significantly, leading to increases on BLUE-1 by $15.84-8.08=7.76$, BLUE-2 by $3.97-1.89=2.08$, Rouge-L by $13.70-6.71=6.99$, IS by $25.56-9.19=16.37$ and FID by $108.50-59.63=48.87$, respectively.
-For the PhotoChat dataset, our end-to-end model still achieves better performances compared with naive Divter, leading improvements of text response generation on BLEU-1 by $12.08-6.28=5.80$, BLEU-2 by $2.94-1.51=1.43$ and Rouge-L by $11.00-5.40=5.60$, and image response Generation on IS by $14.47-4.9=9.57$ and FID by $262.09-75.88=186.21$, respectively.
-Our end-to-end model reaches state-of-the-art performance compared with pipeline models, which verifies the effectiveness of our proposed dialogue system.
-Third, we look at the image response generation scores.
-The best results of IS and FID are $15.8$ and $29.16$, which is from Divter.
-An important reason could be that Sun et al.
-(2022) employ extra image caption resources to train the model, including YFCC100M (Thomee et al., 2016) and ImageNet (Deng et al., 2009).
-Sampling data from the two datasets might significantly influence the scores of IS and FID, because they are computed by the neural network trained on datasets with similar data distribution.
-4. Conclusion
+Figure 4: Image response generation performance for different temperatures of gumbel-softmax.
+The temperature of gumbel-softmax plays an important role in our end-to-end photo-sharing multi-modal dialogue generation model.
+We conduct an analysis experiment to evaluate the relation between the temperature and image generation performance.
+The results are demonstrated in Figure 4, where the x-axis denotes the temperature of gumbel-softmax and the y-axis denotes IS and FIS on the left and right, respectively.
+When the temperature is 1, gumbel-softmax is equivalent to standard softmax.
+As the temperature reduces lower from 1, the IS scores increase significantly, which verifies the effectiveness of gumbel-softmax for our end-to-end model.
+The performance stops increasing after $e^{-4}$, which could be due to the weakened randomness.
+There is a similar tendency for FID, but it is important to note that lower values are better for FID.
+V Conclusion
 In this paper, we proposed the first end-to-end photo-sharing multi-modal dialogue model, which consists of three modules: text dialogue generator, image perceptron and image generator.
-We exploited a large language model as the backbone of the text dialogue generator.
-For the image perceptron, Q-Former equipped the large language model with the ability to see images in the input end.
-For the image generator, we applied the stable diffusion model to enable the large language model to draw images in the output end.
-Particularly, we proposed the dynamic vocabulary transformation matrix and the straight-through gumbel-softmax technique to align the large language model and stable diffusion model and achieve end-to-end gradient propagation.
-Finally, we conducted experiments on PhotoChat and DialogCC datasets to verify the effectiveness of our proposed end-to-end model.
-Compared with two pipeline models, the end-to-end model gained state-of-the-art performances for photo-sharing multi-modal dialogue generation.
-Acknowledgements.
-This work was supported by the stability support plan for Shenzhen colleges and universities.
+We exploited an LLM as the backbone of the text dialogue generator.
+For the image perceptron, Llama-Vision equipped the LLM with the ability to see images in the input end.
+For the image generator, we applied the stable diffusion model to enable the LLM to draw images in the output end.
+Particularly, we proposed the dynamic vocabulary transformation matrix and the straight-through gumbel-softmax technique to align the LLM and stable diffusion model and achieve end-to-end gradient propagation.
+Finally, we conducted experiments on PhotoChat and DialogCC datasets to verify the effectiveness of our end-to-end model.
+Compared with pipeline models, the end-to-end model gained state-of-the-art performances for photo-sharing multi-modal dialogue generation.